OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3

OpenStudy (anonymous):

a line perpendicular to 2x-y+2=0 will be of the form x+2y+k=0 in general , a line perpendicular to ax+by+c=0 will be of the form bx-ay+k=0 ...knowing the point , K can be found and substituted :)

Directrix (directrix):

Perpendicular lines have slopes that multiply to -1.

Directrix (directrix):

y = 2x + 2 y = mx + b is the general equation of a line in slope-intercept form.

Directrix (directrix):

The slope of y = 2x + 2 is 2. What times 2 = -1? @tristynhart

Directrix (directrix):

I have to have some input from you, or I cannot help you knock out this problem.

Directrix (directrix):

@tristynhart Lemme know when you are back on this thread.